OpenStudy (anonymous):

HELP WILL FAN AND MEDAL Solve each equation using factoring. 9x^3 = 25x

OpenStudy (anonymous):

= x( 9x^2 - 25) = x [ ( 3x )^2 - ( 5 )^2 ] = x ( 3x - 5 ) ( 3x + 5 )

OpenStudy (anonymous):

What next?

OpenStudy (anonymous):

that's it, it's completely factored

OpenStudy (anonymous):

Oh, but whats the answer?